OpenStudy (anonymous):

if loga(3)=5 what is the value of a^5

OpenStudy (anonymous):

3

OpenStudy (anonymous):

if you transform that logarithm equation into exponentials, you get a^5 =3

OpenStudy (anonymous):

a^3 = 5 then a = cube root of 5 = 5^1/3 and a^5 = (5^1/3)^5 that is 5^(5/3)

myininaya (myininaya):

\[\log_a(3)=5\] \[a^{\log_a(3)}=a^5\] \[3=a^5\]

OpenStudy (anonymous):

yes a ^5 = 3